OpenStudy (presidentdingy23466):

How do you solve 30(1.2)^x=45+3x.

OpenStudy (daniel.ohearn1):

Use the Lambert W function.

OpenStudy (presidentdingy23466):

what is that

OpenStudy (xapproachesinfinity):

hey there:)

OpenStudy (xapproachesinfinity):

have you started anything or completely stuck

OpenStudy (presidentdingy23466):

haven't started

OpenStudy (presidentdingy23466):

dont know what to do first

OpenStudy (xapproachesinfinity):

hmm what do you know that can simplify this?

OpenStudy (xapproachesinfinity):

I was thinking first hand to make little clean \(30(1.2)^x-3x=45 \Longrightarrow 10(1.2)^x-x=15\)

OpenStudy (xapproachesinfinity):

though it does not seem to break to something easier to solve this we need some other analytical methods

OpenStudy (xapproachesinfinity):

IDK if you learned something of this sort though

OpenStudy (xapproachesinfinity):

algebra alone won't do

OpenStudy (presidentdingy23466):

I'm smart I can probably understand it

OpenStudy (xapproachesinfinity):

in what course are you given this lol?

OpenStudy (presidentdingy23466):

algebra 1

OpenStudy (daniel.ohearn1):

You'll need real analysis for this problem it's post calculus usually.

OpenStudy (zzr0ck3r):

this is a bad question then

OpenStudy (zzr0ck3r):

are you sure there was an x in the exponent and one that is not?

OpenStudy (presidentdingy23466):

yes

OpenStudy (zzr0ck3r):

calculus wont do this either

OpenStudy (xapproachesinfinity):

this is not for alg1 for sure

OpenStudy (daniel.ohearn1):

This foe is beyond you

OpenStudy (daniel.ohearn1):

run

OpenStudy (zzr0ck3r):

I have a masters in math and I have never studied the things needed to solve stuff like this. There is a thing called "analytic continuation of the product log function"

OpenStudy (xapproachesinfinity):

\(\color{blue}{\text{Originally Posted by}}\) @daniel.ohearn1 Use the Lambert W function. \(\color{blue}{\text{End of Quote}}\) I was wondering that this function is about

OpenStudy (xapproachesinfinity):

what*

OpenStudy (zzr0ck3r):

unless of course this is a special case and there is some trivial solution we are not seeing. But in general....this is not undergraduate math (I don't think).

OpenStudy (presidentdingy23466):

nevermind

OpenStudy (xapproachesinfinity):

@zzr0ck3r I'm under gad math so this cannot be solved with complex?

OpenStudy (zzr0ck3r):

I am not sure how this is solved.

OpenStudy (presidentdingy23466):

i think i have the wrong information

OpenStudy (xapproachesinfinity):

oh ok!

OpenStudy (xapproachesinfinity):

could be a trivial solution like you said?

OpenStudy (zzr0ck3r):

well in general \(a^x=x^b\) is not an easy thing to solve, but some are easy \(2^x=x^2\) has an easy solution 2.

OpenStudy (zzr0ck3r):

that is what I meant by trivial, but not in this case.

OpenStudy (zzr0ck3r):

Now I wish I took complex analysis.

OpenStudy (xapproachesinfinity):

yeah seems so, thanks for the clarification :) I just checked the graph they are two real solutions

OpenStudy (xapproachesinfinity):

isn't it a requirement for a math degree?

OpenStudy (zzr0ck3r):

there is a solution for every integer

OpenStudy (zzr0ck3r):

it is not at my school. You just need to take two years of graduate math. I did abstract algebra, topology, graph theory, and a little convex analysis.

OpenStudy (daniel.ohearn1):

No this is not a difficult problem.... Now that I think about it, although it can be solved I imagine with methods of real analysis. For algebra class I would try setting (1.2)^x= (x+15)/10 using your knowledge of graphs and the way they are shaped. At the point of there intersection you will have the solution for x.

OpenStudy (daniel.ohearn1):

There are two intersections amongst the reals.